OpenStudy (anonymous):

Determine the domain and zeros (if any) of the rational function g(x) (x^3-64)/(x^2+5). Also find all intercepts, asymptotes and holes if any

ganeshie8 (ganeshie8):

\(\large g(x) = \dfrac{x^3-64}{x^2+5}\)

ganeshie8 (ganeshie8):

Are there any "x" values that can make denominator become "0" ?

OpenStudy (anonymous):

No

ganeshie8 (ganeshie8):

So, No excluded values for this function that means the Domain the "all real numbers"

OpenStudy (anonymous):

I did get that, I'm stuck on the intercepts, asymptotes and holes.

ganeshie8 (ganeshie8):

To get the zeroes, set the function equal to 0

ganeshie8 (ganeshie8):

For x-intercepts : \(\large g(x) = \dfrac{x^3-64}{x^2+5} = 0\) \(\large \implies x^3-64 = 0\)

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

X=4 that is my x intercept correct.

ganeshie8 (ganeshie8):

Yup !

ganeshie8 (ganeshie8):

knw how to find y-intercepts ?

OpenStudy (anonymous):

No holes, and my ha would be 3/2 and there would be no va.

OpenStudy (anonymous):

Y intercept is 0

OpenStudy (anonymous):

Oops ha is x=1

ganeshie8 (ganeshie8):

there will not be any ha

OpenStudy (anonymous):

No va or ha asymptotes, zero is 4 and intercepts are 4,0

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

"zeroes" is same as "x-intercepts", so ... zeroes = 4 x-intercepts = 4

ganeshie8 (ganeshie8):

how did u get 0 ?

ganeshie8 (ganeshie8):

oh do u mean x-intercept is (4,0) ?

ganeshie8 (ganeshie8):

btw, u still need to find y-intercept

OpenStudy (anonymous):

Not thinking, if I set x to 0 I would get: (0-4)(0+ 4(0)+16) -4(16)=-64 Y=-64?

ganeshie8 (ganeshie8):

hmm doesnt look right

ganeshie8 (ganeshie8):

To get y-intercepts : \(\large g(0) = \dfrac{0^3-64}{0^2+5} \) \(\large g(0) = \dfrac{-64}{5} \)

ganeshie8 (ganeshie8):

\(\large\dfrac{-64}{5} \) is ur y-intercept, right ?

OpenStudy (anonymous):

Yep. Thank you again. Think I am calling it quits for the day

ganeshie8 (ganeshie8):

hold up, are you taight slant/oblique asymptotes yet ?

ganeshie8 (ganeshie8):

if u heard of them before, then u will have to find them as well :)

OpenStudy (anonymous):

???

ganeshie8 (ganeshie8):

nvm

OpenStudy (anonymous):

Ok, have a great Easter!

ganeshie8 (ganeshie8):

you too :)